dukedadon:

The sets of numbers 7, 24, 25 and 9, 40, 41 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples!

dukedadon:

@tranquility care to help?

axie:

Ok so Pythagrean Theorem is Leg^2 x Leg^2 = Hypotonouse^2

axie:

so like 7^2 x 24^2 = H^2

axie:

"explain or show why they are Pythagorean triples." is what we are asked correct?

axie:

"A Pythagorean triple consists of three positive integers a, b, and c, such that a2 + b2 = c2. Such a triple is commonly written (a, b, c), and a well-known example is (3, 4, 5). If (a, b, c) is a Pythagorean triple" Kahn Academy

axie:

So how does the set 7, 24,25 represent this?